On a summer day, the probability that Anne stays outside for more than 3 hours is 80%. If she stays outside for more than 3 hours, the probability that she gets a sunburn is 65%. If she stays outside for less than 3 hours, the probability that she gets a sunburn is 25%.

(a) Create a sample space for this situation showing all the possibilities.

(b) Anne thinks that the probability of getting a sunburn on a summer day is 60%. Is she correct? Explain your reasoning using probabilities.
Answer #1

let D stay outside for more than 3 hours. Getr a sunburn. then from giren data: P (D) - 0.8 PCD)-1-P1D) = 1-0.8 -0.2 P(Elo -0summen :: PC Getting sunburn on day (B) = P( B DD) + P (DNB ple) = P[END] + pleno) 6.52 to.os... Z 0.57 60 y day is PCE). b]Thank you.